OpenStudy (anonymous):

which of the following has a graph that passes through a point (3,1) a. 2x=y-7 b.y=4x-9 c.y=2/3x-5 d.3x-7y=2

OpenStudy (anonymous):

i think i solved this one lol dont post again

OpenStudy (anonymous):

an easy way would be to plug the point in the given options . the equation that will satisfy this point is your answer let me try one put x= 3 and y=1 in the first option (a) 2x=y-7 2(3)=1-7 6=-6 impossible because mathematically 6 is not equal to -6 similarly put x= 3 and y =1 in second option (b) .y=4x-9 1=4(3)-9 1=12-9 1=3 again its mathematically incorrect . can you try this for option c and ?? the one in which right hand side will be equal to left hand side is your answer

OpenStudy (anonymous):

yes ! thank u

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

did you find the correct one ?

OpenStudy (anonymous):

is it d.?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

thanks !